OpenStudy (anonymous):

A recipe for cookies calls for a quarter cup if milk for every 12 cookies. If we want to make 30 cookies, how much milk are we going to need? a.) 1 cup b.) 1/4 cup c.) 2 cups d.) 5/8 cup e.) 1/16 cup I know the answer is D but how did they get the answer? PS. I have problems with my internet so if I disconnect I'll be trying to connect again as fast as I can. Sorry if I make anyone wait.

OpenStudy (kropot72):

The quantity of milk needed for each cookie is found by dividing a quarter by 12: \[amount\ per\ cookie=\frac{\frac{1}{4}}{12}\ cup\] Amount of milk for 30 cookies is \[\frac{\frac{1}{4}}{12} \times 30=\frac{30}{4\times12}=\frac{30}{48}\ cup\] Now you just need to simplify the fraction 30/48

OpenStudy (anonymous):

oh thanks this helps a lot :)

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

thanks!!